# Gaussian consensus processes and their Lyapunov exponents

### v1

@@ Abstract
We introduce a simple dynamic model of opinion formation, in which a finite population of individuals hold vector-valued opinions. At each time step, each individual’s opinion moves towards the mean opinion but is then perturbed independently by a centred multivariate Gaussian random variable, with covariance proportional to the covariance matrix of the opinions of the population. We establish precise necessary and sufficient conditions on the parameters of the model, under which all opinions converge to a common limiting value. Asymptotically perfect correlation emerges between opinions on different topics. Our results are rigorous and based on properties of the partial products of an i.i.d. sequence of random matrices. Each matrix is a fixed linear combination of the identity matrix and a real Ginibre matrix. We derive an analytic expression for the maximal Lyapunov exponent of this product sequence. We also analyze a continuous-time analogue of our model.
@@ 1 Introduction
In this paper, we introduce two simple stochastic processes which we call Gaussian consensus processes. Model A is a discrete-time Markov process and Model B is a continuous-time Markov diffusion. They are fairly reasonable and yet highly mathematically tractable stochastic models of opinion dynamics.
@@
Opinion dynamics is a popular subject in the current mathematical modelling and mathematical social science literature, for example, see Hassani et al. [11], Peralta et al. [22] and references therein. Often the focus is on explanatory power; for example the recent survey by Devia et al. [8] compares the qualitative features of the output of a variety of opinion formation models with opinion data from a selection of real world surveys. In contrast, our focus will be on the theoretical probabilistic properties of a stochastic opinion formation model that is designed to be highly tractable, but nevertheless displays qualitative features which match at least one familiar aspect of opinion dynamics in the real world.
@@
Let us mention a few stochastic opinion dynamics models which are somewhat close to the current paper. In the Weisbuch-Deffuant model [7], $N$ individuals each hold an opinion in the interval $[0,1]$; at each time step two individuals are chosen uniformly at random, and if their opinions are less than $\epsilon$ apart then they each move their opinion towards the mean of their two opinions. The model converges almost surely to a limit in which clusters of individuals all agree perfectly within each cluster, and the clusters are separated by distance at least $\epsilon$. In a noisy variation considered by Piñeda et al. [23], each individual occasionally resamples their opinion from the uniform distribution on $[0,1]$. The bounded confidence cutoff, controlled by the parameter $\epsilon$, is a nonlinearity which induces clustering of the opinions into multiple clusters when the non-linearity is strong enough in comparison to the noise.
@@
In the well-known DeGroot model, individuals correspond to vertices of a graph. Edges correspond to pairs of individuals who influence each other. At each time step, each individual deterministically updates their $\mathbb{R}$-valued opinion to the mean of their neighbours’ opinions at the previous time step. Stern and Livan [24] recently studied a stochastic variant of the DeGroot model in which the opinions are further modified at each time step by the addition of Gaussian noise. Stern and Livan’s equation (3.4) allows the variance of the noise terms to depend on the current state in a general way, which includes the one-dimensional case of our Model A as a special case. However, their main object of study is a model in which the variance of the noise term is constant. In the alternative models considered in Section 7 of [24], the variance of the noise term varies from individual to individual according to how far their opinion is from the empirical mean. Stern and Livan’s main focus is a numerical study of the effect of the community structure of the graph which governs the interactions. In contrast, in our models, the opinions are in $\mathbb{R}^{d}$ and the covariance of the noise term is a function of the empirical covariance of the opinions, while the underlying graph is the complete graph on $N$ vertices.
@@ 1.1 Models
The models that we introduce and study in this paper are mean field (the underlying graph is the complete graph) and linear (there is no bounded confidence cutoff). This makes them probably less realistic as direct models of opinion dynamics in the real world than some of the models mentioned above [^3] Our models could potentially arise as linearizations when studying the stability of a fixed point of a deterministic nonlinear opinion dynamics model when only a small amount of stochastic noise is added; however we will treat them as interesting stochastic processes in their own right..
The novel feature of our models is the combination of Gaussian noise and invariance under affine linear changes of coordinates; these two properties allow us to compute precisely the critical parameters on the boundary between stability and instability, using the classical theory of Lyapunov exponents of random matrix products.
In Model A, $N\geq 2$ individuals each hold real-valued opinions on $d$ topics. Each individual updates their opinions once every day. The opinions of individual $i$ on day $t$ are represented by a row vector $X_{i}(t)\in\mathbb{R}^{d}$. We collect these row vectors in an $N\times d$ matrix $X(t)$ whose $i^{th}$ row is $X_{i}(t)$. The initial condition $X(0)$ may be deterministic or random. For each time $t\in\mathbb{N}_{0}$, we denote by $\mu^{X}(t)$ and $\mathrm{Cov}^{X}(t)$ the empirical mean and empirical covariance matrix of the $N$ opinion vectors at time $t$. The model has two parameters: $\alpha>0$ and $\beta\in[0,1]$. We may think of $\alpha$ as a measure of how much the individuals like to stand out from the crowd, and $\beta$ as a measure of their self-confidence. For each $t\geq 1$, conditional on $X(t-1)$, let $Y_{1}(t),\dots,Y_{N}(t)$ be independent samples from the multivariate Gaussian distribution with mean $\mu^{X}(t-1)$ and $d$ by $d$ covariance matrix $\alpha\,\mathrm{Cov}^{X}(t-1)$. Then for each $1\leq i\leq N$, let
@@
Note that conditional on $X(t-1)$ the $N$ opinion vectors $X_{1}(t),\dots,X_{N}(t)$ are independent. Also, note that $\left(X(t)\right)_{t\in\mathbb{N}_{0}}$ is a discrete-time Markov process.
We will show in Proposition 1 that $\left(\mathrm{Cov}^{X}(t)\right)_{t\geq 0}$ is a Markov chain on its own, and conditional on its entire evolution, the empirical mean process $\left(\mu^{X}(t)\right)_{t\geq 0}$ is a martingale with independent Gaussian increments. This structure enables us to reduce questions about the model to questions about the sequence of partial products of a sequence $M(1),M(2),\dots$ of i.i.d. random $(N-1)\times(N-1)$ matrices. Each matrix $M(t)$ is distributed as a linear combination of the identity and a real Ginibre matrix. A random $k\times k$ matrix is called a real Ginibre matrix if it has i.i.d. standard normal entries, and its distribution is called the real Ginibre ensemble of rank $k$. A precise description of $M(t)$ will be given below in §1.2.
An important question about any opinion dynamics model is whether the opinions converge almost surely to a consensus, i.e. to a finite limit in the opinion space $\mathbb{R}^{d}$. Theorem 1 answers this question for all values of $N,\alpha$, and $\beta$. Our analysis reduces the problem to determining the sign of the maximal Lyapunov exponent of the matrix product sequence. To keep the paper self-contained, we include a brief summary of the key definitions and theorems about Lyapunov exponents of i.i.d. random matrix products. Extending a classical result of Newman [20] about products of real Ginibre matrices, we use the fact that the law of $M(1)$ is invariant under orthogonal conjugation to compute the maximal Lyapunov exponent of our i.i.d. matrix product exactly. For even $N\geq 4$ we express it as an infinite series involving the digamma function, and for odd $N$ we express it in terms of the exponential integral function $\mathrm{Ei}_{1}$. We believe this computation has not previously appeared in the literature. In fact there are only a few random matrix ensembles with nontrivial maximal Lyapunov exponents that are known to be expressible without reference to an inexplicit invariant measure on a projective space; see for example [9, 14, 19, 20]. Our computation of the maximal Lyapunov exponent determines the precise region in parameter space where consensus is reached almost surely. We show that outside of this region the opinion vectors almost surely diverge as $t\to\infty$. In the critical case, on the boundary of the consensus region, almost surely the diameter of the set of opinions is neither bounded nor bounded away from 00 as $t\to\infty$. In the supercritical case the individual opinions and the mean all tend to $\infty$ almost surely.
We also obtain an expression for all $N-1$ Lyapunov exponents, and compute explicit positive lower bounds on the gaps between successive Lyapunov exponents, in terms of the parameters $(N,\alpha,\beta)$. The gap between the first two Lyapunov exponents has an interpretation in the opinion dynamics, given in Theorem 2. The gap is the exponential rate at which the population’s opinions on the $d$ different topics tend towards being perfectly correlated. It is striking that even in a very simple linear model we can observe the formation of a one-dimensional ‘political spectrum’, whose alignment in the $d$-dimensional opinion space is random. This is described in Theorem 3.
Model B is a continuous-time analogue of Model A, with time running over $[0,\infty)$. It is defined in detail in §1.3 below. Briefly, the $N$ opinion vectors diffuse, solving a stochastic differential equation. The diffusion coefficient of each individual’s opinion is defined to be the unique non-negative definite symmetric square root of the empirical covariance matrix of the $N$ opinion vectors in $\mathbb{R}^{d}$. There is also a linear drift towards the empirical mean of the $N$ opinion vectors. We check that the SDE is well-posed using the Araki–Yamagami inequality. We determine the critical strength of the linear drift term, above which consensus is reached almost surely, as a function of $N$. The random matrix product that occurs in the analysis of Model A is replaced in the analysis of Model B by the ellipsoid diffusion process described by Norris, Rogers and Williams [21], who gave an elementary treatment of earlier work by Dynkin and Orihara.
1.2 Results about the discrete time model (Model A)
Recall that $\mu^{X}(t)\in\mathbb{R}^{d}$ denotes the empirical mean opinion at time $t$:
$\mu^{X}(t)=\frac{1}{N}\sum_{k=1}^{N}X_{k}(t)\,.$
The coordinates of $\mu^{X}(t)$ are $\mu_{1}^{X}(t),\mu_{2}^{X}(t),\dots,\mu_{d}^{X}(t)$. Consider the centred $N\times d$ matrix $\bar{X}(t)$ with rows $X_{i}(t)-\mu^{X}(t)$ for $i=1,\dots,N$. Each column of $\bar{X}(t)$ has sum zero. Define
@@ 1.2 Results about the discrete time model (Model A)
Thus $\mathrm{Cov}^{X}(t)$ is the $d$ by $d$ covariance matrix of the empirical distribution at time $t$. Its entry in row $i$ and column $j$ is
@@
We now describe the transition law of the Markov chain $\left(X(t)\right)_{t\geq 0}$. At each time step $t\geq 1$, conditional on $X(t-1)$, $X_{1}(t),\dots,X_{N}(t)$ are independent and for $X_{i}(t)$ has the multivariate Gaussian distribution with mean
$\beta X_{k}(t-1)+(1-\beta)\mu^{X}(t-1)$
@@
$\rho:=\alpha(1-\beta)^{2}\,.$
One can easily check that this description agrees with the one given earlier in the introduction.
@@
Before stating our results, we rule out two special cases. Firstly we assume $\beta\neq 1$, since the case $\beta=1$ corresponds to completely stubborn individuals who never change their opinion. Thus $\beta\in[0,1)$. Secondly, we assume that almost surely there is no nontrivial linear combination of the $d$ opinions on which all $N$ individuals agree. In other words, we assume that $\mathrm{Cov}^{X}(0)$ is positive definite almost surely. For if we start with perfect consensus on some topic represented by a linear functional of the $d$ coordinates, then no individual can ever change their opinion on that topic; we may as well ignore this topic, which has the effect of projecting the model to a lower dimension. As a consequence, we must assume $N\geq d+1$.
@@
We leave it as an open problem whether the conclusion of part (iii) holds in the case where $\lambda_{1}(\alpha,\beta,N)=0$.
Remark 3.
@@
We leave it as an open problem whether the conclusion of part (iii) holds in the case where $\lambda_{1}(\alpha,\beta,N)=0$.
@@
In Proposition 3 in §2.5 we will give closed form analytic expressions for $\lambda_{1}(\alpha,\beta,N)$ in the cases where $N$ is odd.
Figure 1: $\lambda_{1}$ as function of $\alpha,\beta$ for $N=3$
Figure 2: Phase transition of $\lambda_{1}$ for $N=3$. The grey area is where the model is unstable, i.e. $\lambda_{1}\geq 0$.
Figure 3: $\lambda_{1}$ as function of $\alpha,\beta$ for $N=9$
Figure 4: Phase transition of $\lambda_{1}$ for $N=9$
@@
where $\left(G^{(N-1)}\!(t)\right)_{t\geq 0}$ is an i.i.d. sequence drawn from the rank $(N-1)$ real Ginibre ensemble and $I_{N-1}$ is the $(N-1)\times(N-1)$ identity matrix. The quantity $\lambda_{1}$ appearing in Theorem 1 is in fact the largest Lyapunov exponent of the product sequence $\left(P(t)\right)_{t\geq 1}$ defined by
@@
    For parameters $N,\alpha,\beta$, and $\rho=\alpha(1-\beta)^{2}$, for each $k=1,\dots,N-1$ the $k^{th}$ Lyapunov exponent $\lambda_{k}$ of the i.i.d. random matrix product sequence $\left(P(t)\right)_{t\geq 1}$ is given by
    $\lambda_{k}=\lambda_{k}(\alpha,\beta,N)=\frac{1}{2}\mathbb{E}\left(\log\left(\left(\beta W_{k}+\frac{\sqrt{\rho}}{\sqrt{N}}X_{1}\right)^{2}+\sum_{i=2}^{N-k}\frac{\rho}{N}X_{i}^{2}\right)\right)\,,$
    where $W_{k}$ is a random variable distributed as the Euclidean length of the projection of the unit vector $e_{k}$ onto the orthogonal complement of the span of the first $k-1$ rows of the matrix $M(1)$, and $X_{1},\dots,X_{N-k}$ are i.i.d. standard Gaussians, independent of $W_{k}$.
@@
    The Lyapunov spectrum of $\left(P(t)\right)_{t\geq 1}$ is simple, meaning that the $N-1$ Lyapunov exponents $\lambda_{1},\dots,\lambda_{N-1}$ are distinct; moreover
@@
    for every $k$ in the range $21\leq k\leq N-$ 2. (See Theorem 3 for a stronger result about $\lambda_{1}-\lambda_{2}$.)
@@
As $t\to\infty$, $\mathrm{Cov}^{X}(t)/\|\mathrm{Cov}^{X}(t)\|$ converges to a random rank one matrix, regardless of the value of $\lambda_{1}(\alpha,\beta,N)$. Equivalently, there is a random line $\ell$ through 00 in $\mathbb{R}^{d}$ to which the normalized opinions $\hat{X}_{1}(t),\dots,\hat{X}_{N}(t)$ all converge as $t\to\infty$. The exponential rate of convergence is $\lambda_{1}-\lambda_{2}$, which satisfies
@@
as $N\to\infty$ with $\alpha$ and $\beta$ fixed such that $\rho=\alpha(1-\beta)^{2}>0$. If $v_{\ell}$ denotes a unit vector parallel to $\ell$, then $(\hat{X}_{1}(t).v,\dots,\hat{X}_{N}(t).v)$ converges in distribution as $t\to\infty$ to a random variable that is uniformly distributed on the $(N-2)$-dimensional sphere
$S_{N-2}:=\left\{(x_{1},\dots,x_{N})\in\mathbb{R}^{N}\,:\,\sum_{i=1}^{N}x_{i}=0,\;\;\sum_{i=1}^{N}x_{i}^{2}=1\right\}\,.$
@@
$=$ $\left(\frac{\rho^{2}(N-1)}{N^{2}}+2(\beta-1)^{2}\frac{\rho}{N}\right)\left(\mathrm{Cov}_{ii}^{X}(t-1)\,\mathrm{Cov}_{jj}^{X}(t-1)+\mathrm{Cov}^{X}_{ij}(t-1)^{2}\right)\,.$
Fix an $N-1$ by $N$ matrix $V$ whose rows are orthonormal and all orthogonal to the vector $(1,\dots,1)$. Then the second sentence of Proposition 1 may be rephrased by saying that the mapping $X\to\frac{1}{N}X^{T}V^{T}VX$ from the space of real $N\times d$ matrices to the space of non-negative definite symmetric $d$ by $d$ matrices is a strong lumping of the Markov process $\left(X(t)\right)_{t\geq 0}$.

### v2

@@ Abstract
We introduce a simple dynamic model of opinion formation, in which a finite population of individuals hold vector-valued opinions. At each time step, each individual’s opinion moves towards the mean opinion but is then perturbed independently by a centred multivariate Gaussian random variable, with covariance proportional to the covariance matrix of the opinions of the population. We establish precise necessary and sufficient conditions on the parameters of the model, under which all opinions converge to a common limiting value. Asymptotically perfect correlation emerges between opinions on different topics. Our results are rigorous and based on properties of the partial products of an i.i.d. sequence of random matrices. Each matrix is a fixed linear combination of the identity matrix and a real Ginibre matrix. We derive an analytic expression for the maximal Lyapunov exponent of this product sequence. We also analyse a continuous-time analogue of our model.
@@ 1 Introduction
In this paper, we introduce two simple stochastic processes which we call Gaussian consensus processes. Model A is a discrete-time Markov process, and Model B is a continuous-time Markov diffusion. They are fairly reasonable and yet highly mathematically tractable stochastic models of opinion dynamics.
@@
Opinion dynamics is a popular subject in the current mathematical modelling and social science literature, for example, see Hassani et al. [11], Peralta et al. [24] and references therein. Often the focus is on explanatory power; for example, the recent survey by Devia et al. [8] compares the qualitative features of the output of a variety of opinion formation models with opinion data from a selection of real-world surveys. Our focus, however, will be on the theoretical probabilistic properties of a stochastic opinion formation model that is designed to be highly tractable, but nevertheless displays qualitative features which match at least one familiar aspect of opinion dynamics in the real world. Another famous opinion dynamics model, the voter model, concerns discrete opinions evolving on a graph, with individuals being influenced by their immediate neighbours. This is not closely related to the topic of the present paper, because our models are mean field with opinions in a continuum.
@@
Let us mention a few stochastic opinion dynamics models which are somewhat closely related to the current paper. In the Weisbuch-Deffuant model [7], $N$ individuals each hold an opinion in the interval $[0,1]$; at each time step two individuals are chosen uniformly at random, and if their opinions are less than $\epsilon$ apart then they each move their opinion towards the mean of their two opinions. The model converges almost surely to a limit in which clusters of individuals all agree perfectly within each cluster, and the clusters are separated by distance at least $\epsilon$. In a noisy variation considered by Piñeda et al. [25], each individual occasionally resamples their opinion from the uniform distribution on $[0,1]$. The bounded confidence cutoff, controlled by the parameter $\epsilon$, is a nonlinearity which induces clustering of the opinions into multiple clusters when the non-linearity is strong enough in comparison to the noise.
@@
In the well-known DeGroot model, individuals correspond to vertices of a graph. Edges correspond to pairs of individuals who influence each other. At each time step, each individual deterministically updates their $\mathbb{R}$-valued opinion to the mean of their neighbours’ opinions at the previous time step. Stern and Livan [27] recently studied a stochastic variant of the DeGroot model in which the opinions are further modified at each time step by the addition of Gaussian noise. Stern and Livan’s equation (3.4) allows the variance of the noise terms to depend on the current state in a general way, which includes the one-dimensional case of our Model A as a special case. However, their main object of study is a model in which the variance of the noise term is constant. In the alternative models considered in Section 7 of [27], the variance of the noise term varies from individual to individual according to how far their opinion is from the empirical mean. Stern and Livan’s main focus is a numerical study of the effect of the community structure of the graph which governs the interactions. In contrast, in our models, the opinions are in $\mathbb{R}^{d}$, and the covariance of the noise term is a function of the empirical covariance of the opinions, while the underlying graph is the complete graph on $N$ vertices.
Finally, we would like to mention a few very recent papers related to opinion dynamics and somewhat relevant to our paper: [16, 26, 28].
@@ 1.1 Models
The models that we introduce and study in this paper are mean field (the underlying graph is the complete graph) and linear (there is no bounded confidence cutoff). The novel feature of our models is the combination of Gaussian noise and invariance under affine linear changes of coordinates; these two properties allow us to compute precisely the critical parameters on the boundary between stability and instability, using the classical theory of Lyapunov exponents of random matrix products.
In Model A, $N\geq 2$ individuals each hold real-valued opinions on $d$ topics. Each individual updates their opinions once every day. The opinions of individual $i$ on day $t$ are represented by a row vector $X_{i}(t)\in\mathbb{R}^{d}$. We collect these row vectors in an $N\times d$ matrix $X(t)$ whose $i^{th}$ row is $X_{i}(t)$. The initial condition $X(0)$ may be deterministic or random. For each time $t\in\mathbb{N}_{0}$, we denote by $\mu^{X}(t)$ and $\mathrm{Cov}^{X}(t)$ the empirical mean and empirical covariance matrix of the $N$ opinion vectors at time $t$. The model has two parameters: $\alpha>0$ and $\beta\in[0,1]$. We may think of $\alpha$ as a measure of how much the individuals like to stand out from the crowd, and $\beta$ as a measure of their self-confidence. For each $t\geq 1$, conditional on X(t1)X(t-1), let $Y_{1}(t),\dots,Y_{N}(t)$ be independent samples from the multivariate Gaussian distribution with mean $\mu^{X}(t-1)$ and $d\times d$ covariance matrix $\alpha\,\mathrm{Cov}^{X}(t-1)$. Then for each $1\leq i\leq N$, let
@@
Note that conditional on X(t1)X(t-1) the $N$ opinion vectors $X_{1}(t),\dots,X_{N}(t)$ are independent. Also, note that $\left(X(t)\right)_{t\in\mathbb{N}_{0}}$ is a discrete-time Markov process.
Formally,
$\mu^{X}(t)=\frac{1}{N}\sum_{k=1}^{N}X_{k}(t)\,$,
and the coordinates of $\mu^{X}(t)$ are $\mu_{1}^{X}(t),\mu_{2}^{X}(t),\dots,\mu_{d}^{X}(t)$. Consider the centred $N\times d$ matrix $\bar{X}(t)$ with rows $X_{i}(t)-\mu^{X}(t)$ for $i=1,\dots,N$. Each column of $\bar{X}(t)$ has sum zero. Define
@@
Thus $\mathrm{Cov}^{X}(t)$ is the $d\times d$ covariance matrix of the empirical distribution at time $t$. Its entry in row $i$ and column $j$ is
@@
We now describe the transition law of the Markov chain $\left(X(t)\right)_{t\geq 0}$. At each time step $t\geq 1$, conditional on X(t1)X(t-1), the opinion vectors $X_{1}(t),\dots,X_{N}(t)$ are independent and for each $i=1,\dots,N$, $X_{i}(t)$ has the multivariate Gaussian distribution with mean
$\beta X_{i}(t-1)+(1-\beta)\mu^{X}(t-1)$
@@
$\rho:=\alpha(1-\beta)^{2}\,.$ (2)
We will show in Proposition 1 that $\left(\mathrm{Cov}^{X}(t)\right)_{t\geq 0}$ is a Markov chain on its own, and conditional on its entire evolution, the empirical mean process $\left(\mu^{X}(t)\right)_{t\geq 0}$ is a martingale with independent Gaussian increments. This structure enables us to reduce questions about the model to questions about the sequence of partial products of a sequence $M(1),M(2),\dots$ of i.i.d. random $(N-1)\times(N-1)$ matrices. Each matrix $M(t)$ is distributed as a linear combination of the identity and a real Ginibre matrix. A random $k\times k$ matrix is called a real Ginibre matrix if its entries are i.i.d. standard normal random variables. Its distribution is called the real Ginibre ensemble of rank $k$. A precise description of $M(t)$ will be given below in Section 1.2.
An important question about any opinion dynamics model is whether the opinions converge almost surely to a consensus, i.e. to a finite limit in the opinion space $\mathbb{R}^{d}$. Theorem 1 answers this question for all values of $N,\alpha$, and $\beta$. Our analysis reduces the problem to determining the sign of the maximal Lyapunov exponent of the matrix product sequence. To keep the paper self-contained, we include a brief summary of the key definitions and theorems about Lyapunov exponents of i.i.d. random matrix products. Extending a classical result of Newman [22] about products of real Ginibre matrices, we use the fact that the law of $M(1)$ is invariant under orthogonal conjugation to compute the maximal Lyapunov exponent of our i.i.d. matrix product exactly. For even $N\geq 4$ we express it as an infinite series involving the digamma function, and for odd $N$ we express it in terms of the exponential integral function $\mathrm{Ei}_{1}$. We believe this computation has not previously appeared in the literature. In fact there are only a few random matrix ensembles with nontrivial maximal Lyapunov exponents that are known to be expressible without reference to an inexplicit invariant measure on a projective space; see for example [9, 14, 21, 22]. Our computation of the maximal Lyapunov exponent determines the precise region in parameter space where consensus is reached almost surely. We show that outside of this region the opinion vectors almost surely diverge as $t\to\infty$. In the critical case, on the boundary of the consensus region, almost surely the diameter of the set of opinions is neither bounded nor bounded away from 0 as $t\to\infty$. In the supercritical case, the individual opinions and the mean all tend to $\infty$ almost surely.
We also obtain an expression for all N1N-1 Lyapunov exponents, and compute explicit positive lower bounds on the gaps between successive Lyapunov exponents, in terms of the parameters $(N,\alpha,\beta)$. The gap between the first two Lyapunov exponents has an interpretation in the opinion dynamics, given in Theorem 2. The gap is the exponential rate at which the population’s opinions on the $d$ different topics tend towards being perfectly correlated. It is striking that even in a very simple linear model we can observe the formation of a one-dimensional ‘political spectrum’, whose alignment in the $d$-dimensional opinion space is random. This is described in Theorem 3.
Model B is a continuous-time analogue of Model A, with time running over $0,\infty)$. It is motivated by considering the possible scaling limit of Model A as the updates become more frequent and less strong, though we do not prove such a scaling limit result in this paper; we pose it as an open problem to identify the correct scaling regime. It is defined in detail in Section 1.3 below. Briefly, the $N$ opinion vectors diffuse, solving a stochastic differential equation. The diffusion coefficient of each individual’s opinion is defined to be the unique non-negative definite symmetric square root of the empirical covariance matrix of the $N$ opinion vectors in $\mathbb{R}^{d}$. There is also a linear drift towards the empirical mean of the $N$ opinion vectors. We check that the SDE is well-posed using the Araki–Yamagami inequality. We determine the critical strength of the linear drift term, above which consensus is reached almost surely, as a function of $N$. The random matrix product that occurs in the analysis of Model A is replaced in the analysis of Model B by the ellipsoid diffusion process described by Norris, Rogers and Williams [23], who gave an elementary treatment of earlier work by Dynkin and Orihara.
1.2 Results about the discrete time model (Model A)
@@
Before stating our results, we rule out two special cases. Firstly we assume $\beta\neq 1$, since the case $\beta=1$ corresponds to completely stubborn individuals who never change their opinion. Thus $\beta\in[0,1)$. Secondly, we assume that almost surely there is no nontrivial linear combination of the $d$ opinions on which all $N$ individuals agree. In other words, we assume that $\mathrm{Cov}^{X}(0)$ is positive definite almost surely. For if we start with perfect consensus on some ‘mixture of topics’ represented by a linear functional of the $d$ coordinates, then no individual can ever change their opinion on this mixture; we may as well ignore this direction, which has the effect of projecting the model to a lower dimension. As a consequence, we must assume $N\geq d+1$.
@@
Some illustrations of how $\lambda_{1}$ depends on $\alpha$, $\beta$ and $N$, and relevant phase transitions, are given in Figures 1 and 2.
Figure 1: $\lambda_{1}$ as function of $\alpha,\beta$ for $N=3$ (left) and $N=9$ (right).
Figure 2: Phase transition of $\lambda_{1}$ for $N=3$ (left) and $N=9$ (right). The grey area is where the model is unstable, i.e. $\lambda_{1}\geq 0$.
@@
We leave it as an open problem whether the conclusion of part (iii) holds in the case where $\lambda_{1}(\alpha,\beta,N)=0$. Namely, we do not know if the divergence of the opinions occurs in all directions, or rather only in some subspace of a lower dimension.
@@
In Proposition 3 in Section 2.5 we will give closed form analytic expressions for $\lambda_{1}(\alpha,\beta,N)$ in the cases where $N$ is odd.
@@
where $\left(G^{(N-1)}\!(t)\right)_{t\geq 0}$ is an i.i.d. sequence drawn from the rank (N1)(N-1) real Ginibre ensemble and $I_{N-1}$ is the $(N-1)\times(N-1)$ identity matrix. The quantity $\lambda_{1}$ appearing in Theorem 1 is in fact the largest Lyapunov exponent of the product sequence $\left(P(t)\right)_{t\geq 1}$ defined by
@@
Indeed, by Section 2.2.1, the limit ${\lambda_{1}}=\lim_{t\to\infty}\frac{1}{t}\log\|P(t)\|$ exists; Theorem 1 shows $\lambda_{1}<0$ iff Model A reaches consensus a.s., with the critical and supercritical cases treated there.
@@
    For parameters $N,\alpha,\beta$, and $\rho=\alpha(1-\beta)^{2}$, for each $k=1,\dots,N-1$, the $k^{th}$ Lyapunov exponent $\lambda_{k}$ of the i.i.d. random matrix product sequence $\left(P(t)\right)_{t\geq 1}$ is given by
    $\lambda_{k}=\lambda_{k}(\alpha,\beta,N)=\frac{1}{2}\mathbb{E}\left(\log\left(\left(\beta W_{k}+\frac{\sqrt{\rho}}{\sqrt{N}}X_{1}\right)^{2}+\frac{\rho}{N}\sum_{i=2}^{N-k}X_{i}^{2}\right)\right)\,$,
    where $W_{k}$ is a random variable distributed as the Euclidean length of the projection of the unit vector $e_{k}$ onto the orthogonal complement of the span of the first k1k-1 rows of the matrix $M(1)$, and $X_{1},\dots,X_{N-k}$ are i.i.d. standard Gaussians, independent of $W_{k}$.
@@
    The Lyapunov spectrum of $\left(P(t)\right)_{t\geq 1}$ is simple, meaning that the N1N-1 Lyapunov exponents $\lambda_{1},\dots,\lambda_{N-1}$ are distinct; moreover
@@
    for every $k$ in the range $1\leq k\leq N-2$. (See Theorem 3 for a stronger result about $\lambda_{1}-\lambda_{2}$.)
The interpretation of Theorem 2 is that there is a random direction in $\mathbb{R}^{d}$ in which the opinions line up in the limit $t\to\infty$. The law of this direction is the unique invariant measure on $P(\mathbb{R}^{d})$ under the action of the group of linear isometries of the quadratic form $\left(\bar{X}(0)^{T}\,\bar{X}(0)\right)^{-1}$. The convergence of the population of opinions to one-dimensionality occurs exponentially with rate $\lambda_{1}-\lambda_{2}$; later we will show that $\lambda_{1}-\lambda_{2}$ behaves like $1/N$ in the case $\beta=0$, and is at least $c(\alpha,\beta)/N^{2}$ in the general case.
@@
As $t\to\infty$, $\mathrm{Cov}^{X}(t)/\|\mathrm{Cov}^{X}(t)\|$ converges to a random rank one matrix, regardless of the value of $\lambda_{1}(\alpha,\beta,N)$. Equivalently, there is a random line $\ell$ through 0 in $\mathbb{R}^{d}$ to which the normalized opinions $\hat{X}_{1}(t),\dots,\hat{X}_{N}(t)$ all converge a.s. as $t\to\infty$. The exponential rate of convergence is $\lambda_{1}-\lambda_{2}$, which satisfies
@@
as $N\to\infty$ with $\alpha$ and $\beta$ fixed such that $\rho=\alpha(1-\beta)^{2}>0$. If $v$ denotes a unit vector parallel to $\ell$, then $(\hat{X}_{1}(t).v,\dots,\hat{X}_{N}(t).v)$ converges in distribution as $t\to\infty$ to a random variable that is uniformly distributed on the (N2)(N-2)-dimensional sphere
$\mathbb{S}^{N-2}:=\left\{(x_{1},\dots,x_{N})\in\mathbb{R}^{N}\,:\,\sum_{i=1}^{N}x_{i}=0,\;\;\sum_{i=1}^{N}x_{i}^{2}=1\right\}\,.$
It follows from Theorem 3 that for any two topics, i.e. coordinates $1\leq j<k\leq d$, we have asymptotically perfect correlation in the people’s opinions on those topics:
$\mathrm{Corr}(X_{\cdot,j}(t),X_{\cdot,k}(t))\to\pm 1\text{ as $t\to\infty$}.$
The mechanism by which this occurs is that the products of the matrices that update the covariance matrix contract directions in projective space, at asymptotic rate $\lambda_{1}-\lambda_{2}>0$, towards a random direction, yielding the rank one covariance matrix and perfect correlation in the limit.
@@
$\hskip-22.76219pt=\left(\frac{\rho^{2}(N-1)}{N^{2}}+2(\beta-1)^{2}\frac{\rho}{N}\right)\left(\mathrm{Cov}_{ii}^{X}(t-1)\,\mathrm{Cov}_{jj}^{X}(t-1)+\mathrm{Cov}^{X}_{ij}(t-1)^{2}\right)\,.$
Fix an N1N-1 by $N$ matrix $V$ whose rows are orthonormal and all orthogonal to the vector $(1,\dots,1)$. Then the second sentence of Proposition 1 may be rephrased by saying that the mapping $X\to\frac{1}{N}X^{T}V^{T}VX$ from the space of real $N\times d$ matrices to the space of non-negative definite symmetric $d$ by $d$ matrices is a strong lumping of the Markov process $\left(X(t)\right)_{t\geq 0}$.